OpenStudy (anonymous):

percent of change 2.5a to 12.5a

OpenStudy (anonymous):

12.5-2.5/2.5 = 4 which is 400% incease

OpenStudy (anonymous):

www.tdk-lambda.com/products/sps/catalog/eng/vs_e.pdf

OpenStudy (anonymous):

You added ^^ when you should have subtracted and taken the absolute value.

hero (hero):

You're right, but my formula is correct

OpenStudy (anonymous):

No, the formula is actually incorrect as well. http://virtual.mjc.edu/rlamont/aplia%20hints/change.bmp

hero (hero):

\[\% \Delta = \frac{a - b}{a} \times 100 = \frac{|2.5a - 12.5|}{2.5} \times 100 = \frac{|-10|}{2.5} \times 100 = \frac{1000}{25} = 400\]

hero (hero):

Nope, I just proved it is correct

OpenStudy (anonymous):

That's what I had.

hero (hero):

Trust me, the formula is correct. How do you explain getting the same exact answer?

hero (hero):

If used correctly, the formula is correct.

OpenStudy (anonymous):

Just saying that you derived the incorrect answer when you used it before.

hero (hero):

You pointed out my error. I fixed it. Why are you still complaining?

OpenStudy (anonymous):

Yes, I do. My message is that, for the good of us all, you should never engage in an endless round of finger pointing. You should never even try to do such an iniquitous thing. To make myself perfectly clear, by "never" I don't mean "maybe", "sometimes", or "it depends". I mean only that if one could get a Ph.D. in Colonialism, then and only then would you be the first in line to have one.

OpenStudy (anonymous):

QUESTION CLOSED!!!

hero (hero):

@brainshot3 , I have no idea what you are talking about. Who did I point a finger at?

hero (hero):

The only thing I said was that my formula is correct. If I pointed a finger at someone, it was at myself. I even acknowledged that I made a mistake. I don't know what you want from me.

OpenStudy (anonymous):

All opposition. You are trying to constate your belief by continuing to dither on such a trivial topic. Drop it.

hero (hero):

I don't really believe you're in a position to tell me what to do, but okay.